The edges of a cube are each 10 cm. All of the edges of a cube are reassembled to form a square. In square centimeters, compute the difference:
(Area of square) – (Surface area of cube).
A. –200 B. 0 C. 300 D. 400 E. 500

Answer by richwmiller(17219)

This cube has 12 edges of equal size 10cm.
12*10=120
now if we make a square then each side will be 30 cm
30*30=900
the surface area is 6*10*10=600
900-600=300
letter c